OpenStudy (anonymous):

um i have another one @jagr2713 and @sleepyjess do you mind helping me? Last one I promise :)

jagr2713 (jagr2713):

Sure :D

OpenStudy (anonymous):

Awesome

OpenStudy (sleepyjess):

I will let @jagr2713 get this one, I have to go

jagr2713 (jagr2713):

why D:

OpenStudy (anonymous):

ok Here it is \[1/3x +1/3 (2x-5 )=3 1/2\]

OpenStudy (anonymous):

awww ok @sleepyjess

OpenStudy (anonymous):

my mistake its not 31 over 2 its 3 and 1/2

jagr2713 (jagr2713):

oh there we go i was like what?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

and those are fractions not division problems haha

jagr2713 (jagr2713):

\[\frac{ 1 }{ 3 }x+\frac{ 1 }{ 3 }(2x-5)=3\frac{ 1 }{ 2 }\]

jagr2713 (jagr2713):

?

OpenStudy (anonymous):

yea thats it ...how did you do that?

jagr2713 (jagr2713):

press the equations button

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh i figured it out haha nvm

jagr2713 (jagr2713):

me to lol :D

jagr2713 (jagr2713):

wait figure out the question the equation thing?

OpenStudy (anonymous):

the equation button thing

jagr2713 (jagr2713):

oh ok lets finish what were doing

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have no idea what to do on this one

jagr2713 (jagr2713):

ok first we convert the mix numbers\[3\frac{ 1 }{ 3 }\] into an improper fraction first by multiplying the denominator (2) by the whole number part (3) and add the numerator (1) to get the new numerator. Place the new numerator (7) over the old denominator (2). and we get \[\frac{ 1 }{ 3 }x+\frac{ 1 }{ 3}(2x-5)=\frac{ 7 }{ 5 }\]

jagr2713 (jagr2713):

then we simply each term and what do we get?

OpenStudy (anonymous):

um i have no idea i hate fraction equations

jagr2713 (jagr2713):

so we multiply 1/3 by x to get (1/3)x \[(\frac{ 1 }{ 3 })x+\frac{ 1 }{ 3}(2x-5)=\frac{ 7 }{ 2 }\]

jagr2713 (jagr2713):

then we simplify \[\frac{ x }{ 3 }+\frac{ 1 }{ 3 }(2x-5)=\frac{ 7 }{ 2 }\]

jagr2713 (jagr2713):

ok what do we multiply next?

OpenStudy (anonymous):

-_- i don't know :( um inside the () ?

jagr2713 (jagr2713):

close but we multiply 1/3 by 2x-5 to get (1/3)(2x-5) \[\frac{ x }{ 3 }+(\frac{ 1 }{ 3 })(2x-5)=\frac{ 7 }{ 2 }\]

jagr2713 (jagr2713):

then we apply the distributive property \[\frac{ x }{ 3 }+(\frac{ 1 }{ 3 })(2x)+(\frac{ 1 }{ 3 })(-5)=\frac{ 7 }{ 2 }\]

jagr2713 (jagr2713):

what do you think we do with 1/3 and 2x

OpenStudy (anonymous):

make it 2 and 1 /3 x?

jagr2713 (jagr2713):

close again but we simplify to make it \[\frac{ x }{ 3 }+\frac{ 2x }{ 3 }+(\frac{ 1 }{ 3 })(-5)=\frac{ 7 }{ 2 }\]

jagr2713 (jagr2713):

from reading that what do we do to 1/3 and -5

OpenStudy (anonymous):

multiply

jagr2713 (jagr2713):

simplify yes and what do we get

OpenStudy (anonymous):

um -1.67

jagr2713 (jagr2713):

now like the other one we do we get \[\frac{ x }{ 3 }+\frac{ 2x }{ 3 }+\frac{ -5 }{ 3 }=\frac{ 7 }{ 2 }\]

jagr2713 (jagr2713):

next we move the negative in the front of the fraction, what do we get

OpenStudy (anonymous):

x/3 + -5/3 + 2x/3 = 7/2

jagr2713 (jagr2713):

brb

OpenStudy (anonymous):

k

jagr2713 (jagr2713):

correct but you just flip 2x/3, -5/3

jagr2713 (jagr2713):

\[\frac{ x }{ 3 }+\frac{ 2x }{ 3 }-\frac{ 5 }{ 3 }=\frac{ 7 }{ 2 }\]

OpenStudy (anonymous):

i have to go in a min...can we kinda speed this up? not meaning to be rude

jagr2713 (jagr2713):

ok now for me i usually combine the to numerators u can if u want but i am going to combine and get \[\frac{ x }{3 }+\frac{ 2x-5 }{ 3 }=\frac{ 7 }{ 2 }\] the den stays the same

jagr2713 (jagr2713):

umm sure

OpenStudy (anonymous):

:)

jagr2713 (jagr2713):

ok but u have at least 5mins

OpenStudy (anonymous):

no maybe 2

OpenStudy (anonymous):

2 mins that is

jagr2713 (jagr2713):

ok

jagr2713 (jagr2713):

Combine x/3+(2 x-5)/3 into a single fraction. x/3+(2 x-5)/3 = ((2 x-5)+x)/3: (2 x+x-5)/3 = 7/2 Group like terms in 2 x+x-5. Grouping like terms, 2 x+x-5 = (x+2 x)-5: ((x+2 x)-5)/3 = 7/2 Add like terms in x+2 x. x+2 x = 3 x: (3 x-5)/3 = 7/2

OpenStudy (anonymous):

whoa

jagr2713 (jagr2713):

u want me to go fast right?

OpenStudy (anonymous):

yes haha

jagr2713 (jagr2713):

Multiply both sides by a constant to simplify the equation. Multiply both sides of (3 x-5)/3 = 7/2 by 3: (3 (3 x-5))/3 = 3×7/2 Express 3×7/2 as a single fraction. 3×7/2 = (3×7)/2: (3 (3 x-5))/3 = (3×7)/2 Cancel common terms in the numerator and denominator of (3 (3 x-5))/3. (3 (3 x-5))/3 = 3/3×(3 x-5) = 3 x-5: 3 x-5 = (3×7)/2 Multiply 3 and 7 together. 3×7 = 21: 3 x-5 = 21/2 Isolate terms with x to the left hand side. Add 5 to both sides: 3 x+(5-5) = 21/2+5 Look for two terms that sum to zero. 5-5 = 0: 3 x = 21/2+5 Put the fractions in 21/2+5 over a common denominator. Put 21/2+5 over the common denominator 2. 21/2+5 = 21/2+(2×5)/2: 3 x = 21/2+(2×5)/2 Multiply 2 and 5 together. 2×5 = 10: 3 x = 21/2+10/2 Add the fractions over a common denominator to a single fraction. 21/2+10/2 = (21+10)/2: 3 x = (21+10)/2

jagr2713 (jagr2713):

whats next?

OpenStudy (anonymous):

wow 0.0

jagr2713 (jagr2713):

you dont get it?

OpenStudy (anonymous):

not really

jagr2713 (jagr2713):

yea i cant do it in 2mins lol

OpenStudy (anonymous):

well can you just give me the answer? please?

jagr2713 (jagr2713):

i cant but i am almost done :D just couple more secs

OpenStudy (anonymous):

:( fine

jagr2713 (jagr2713):

ok now we evaluate 21+10 so we get 31 \[3x=\frac{ 31 }{ 2 }\]

jagr2713 (jagr2713):

now we solve for x so we divide both sides by 3 \[x=\frac{ 31 }{ 2*3 }\]

jagr2713 (jagr2713):

whats the answer?

jagr2713 (jagr2713):

you there @koolkat13

OpenStudy (anonymous):

5.16

jagr2713 (jagr2713):

no this is the easiest part what is 3*2

OpenStudy (anonymous):

6

jagr2713 (jagr2713):

so the answer is?

OpenStudy (anonymous):

31/6 = 5.167

jagr2713 (jagr2713):

31/6 is correct but you dont have to divide anymore b.c 31 doesnt go into 6 so leave as it is ;D

OpenStudy (anonymous):

oh ok hehe

jagr2713 (jagr2713):

:D now you may leave

OpenStudy (anonymous):

haha ok thx bye :D

jagr2713 (jagr2713):

bye